OpenStudy (anonymous):

what is the simplest form of the expression x^2yz/y^2

OpenStudy (anonymous):

Do you mean:\[\frac{ x^2yz }{ y^2 }\]

OpenStudy (anonymous):

Is there anything that you see that would 'cancel' out with eachother?

OpenStudy (phi):

do you know that \(y^2\) is short for \( y \cdot y\) ? do you know that \[ \frac{y}{ y \cdot y}= \frac{\cancel{y}}{\cancel{y}\cdot y}= \frac{1}{y}\]

OpenStudy (phi):

that works because y/y (or anything divided by itself) is 1. 1 times anything is the anything.

OpenStudy (phi):

can you do your problem ?